# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v2

@@ Abstract
Swarms of autonomous interactive drones, with the support of recharging technology, can provide compelling sensing capabilities in Smart Cities, such as traffic monitoring and disaster response. Existing approaches, including distributed optimization and deep reinforcement learning (DRL), aim to coordinate drones to achieve cost-effective, high-quality navigation, sensing, and charging. However, they face grand challenges: short-term optimization is not effective in dynamic environments with unanticipated changes, while long-term learning lacks scalability, resilience, and flexibility. To bridge this gap, this paper introduces a new progressive approach that combines short-term plan generation and selection based on distributed optimization with a DRL-based long-term strategic scheduling of flying direction. Extensive experimentation with datasets generated from realistic urban mobility underscores an outstanding performance of the proposed solution compared to state-of-the-art. We also provide compelling new insights about the role of drones density in different sensing missions, the energy safety of drone operations and how to prioritize investments for key locations of charging infrastructure.
@@ I Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities [1]. With their mobility, autonomy, diverse sensors and the unique vantage view, drones have been widely used in several applications, such as reporting traffic violations, congestion and accidents at early stage [2], as well as mapping natural disasters [3]. Furthermore, the inherent limitations in battery capacity of drones influence their spatio-temporal coverage. Drones require recharging during long operations with supporting facilities (e.g., charging stations). Recently, scholars have introduced the wireless charging and commercially available charging platforms for long-term sensing [4].
@@
Inspired by these studies, this work tackles the problem of coordinating swarms of drones for improved navigation, sensing and recharging. The goal is to enhance sensing quality in a large-scale area using both drones and charging stations while minimizing the overall involved costs. This endeavor requires highly adaptive and optimized drone operations. Previous work [5] employs distributed optimization methods for efficient drones path planning and task assignment to enhance sensing quality. These methods enable drones to autonomously self-organize and self-assign their tasks collectively, while adapting to a dynamic environment. However, these methods are reactive to the current state of the environment rather than proactive to address long-term operational challenges stategically. The navigation and sensing of drones are significantly influenced by two factors: their flying directions and future requirements for sensor data on the map. For example, if drones have information about the expected increase in traffic flow in certain areas, even if those areas are currently sparsely populated with vehicles, they can strategically choose to fly to those locations at early stage. This strategic decision can be more effective for tasks such as vehicle detection, although in short-term the benefits are not evident, exhibiting in this way a "slower is faster" effect: instead of prioritizing quick gains, a long-term planning approach is required to optimize the overall effectiveness and efficiency of drone sensing operations.
@@
Deep reinforcement learning (DRL) tackles the challenge of long-term optimization. It accounts for long-term and iterative actions through the Bellman equation, which considers the discounted cumulative reward of agents’ actions [6]. Drones, as heterogeneous agents with distinct behavior patterns and action spaces, can benefit from the powerful deep neural networks employed in DRL, enabling an effective handling of sophisticated state space and time-varying environments. Nevertheless, the curse of dimensionality restricts drones’ sensing capabilities to small-scale applications as the number of drones and time periods increase. One approach to mitigate this limited scalability is to use a centralized model [7], which performs collaborative scheduling on the actions of all agents. However, this approach is centralized and does not provide drones with privacy, resilience and autonomy to self-assign their sensing tasks. As a result, drones need to become more flexible and adaptive to real-time environmental changes without compromising system resilience, against single points of failure.
@@
Given the aforementioned challenges, there is an opportunity for distributed optimization and DRL to leverage each other’s strengths and mitigate their respective limitations. Therefore, in this paper, we design a new framework for the coordination of swarms of drones, Distributed Optimization and (deep) Reinforcement Learning (DO-RL). This framework involves a novel multi-agent DRL algorithm that builds upon on a collective learning approach known for its remarkable performance in terms of scalability (support of a large number of software agents), efficiency (low communication and computational cost), decentralization and resilience [8, 9, 10, 11, 12]. DO-RL relies on a learning model to determine the drones’ strategic navigation and flying directions for sensing, which also determines the recharging stations, while leaving the planning and selection of specific sensing operations to distributed optimization, i.e., collective learning [8, 9]. This approach significantly reduces the computational complexity of agents, facilitates efficient information sharing, maintains the privacy and autonomy of each agent to independently decide its navigation and sensing actions. Furthermore, a set of auxiliary approaches is introduced, including a plan generation strategy that provides high-quality and diverse options for drones navigation and sensing, as well as a periodic state update. To evaluate the effectiveness of DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
@@
The contributions of this paper are outlined as follows: (i) The first attempt to solve the coordination problems of distributed sensing, recharging and navigation by a swarm of autonomous interactive drones. (ii) A novel approach to address this problem, DO-RL, by combining distributed optimization with deep reinforcement learning. (iii) The applicability of decentralized planning and selection of drone operations using collective learning [8, 9] to improve scalability, resilience and flexibility. (iv) A novel DRL algorithm for drones to optimize their navigation, data collection and recharging. (v) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (vi) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (vii) An open-source implementation of the proposed approach and an open dataset [13] containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
@@ II Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as a combinatorial optimization problem, resembling the Traveling Salesmen Problem [5, 21]. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The digraph-based methods, such as greedy approach, [22], ant colony optimization [23], and genetic algorithm [24], are introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, inspection delay, events detection rate and the cost of flying trajectories. However, these works lack comprehensive evaluation in complex sensing scenarios, such as varying time periods, charging stations, and different sensing distributions. Moreover, these centralized methods suffer from a risk of single points of failure [25], wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms.
@@
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective when addressing the distributed sensing problem with drones using complex optimization objectives, limited energy consumption, and a large number of heterogeneous agents. Bai et al. [27] proposed a distributed DRL-based algorithm for UAV swarm flocking control that demonstrates robust performance in unknown environments, effectively addressing universal challenges such as limited perception capability, communication delay, and collision avoidance. Ding et al. [17] integrated the mobile crowdsensing (i.e., human equipped with mobile sensing devices such as smartphones) into UAV sensing, which overcomes the limitations of both human mobility and drone battery capacity. Nevertheless, these centralized approaches lack privacy and autonomy for drones in self-assignment tasks, reducing flexibility, adaptability to the real-time environmental changes and increasing vulnerability to single points of failure. Omoniwa et al. [18] presented a decentralized approach that leverages DRL and shared information from neighbors to improve the energy efficiency of drones in providing wireless connectivity to ground users in dynamic environments. However, it solves a different problem confined to a small number of drones and time periods.
Recharging assignment. Ealier work [28, 29] has studied the recharging scheduling problem: drones efficiently collect sensor data while charging from stations. However, scalability becomes a challenge since the algorithms used in these works only deal with a small number of drones. Zhao et al. [19] proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces, specifically for detecting available parking spots. Their solution incorporates the task selection and scheduling using DRL, and adaptive charging scheduling. However, these methods are centralized and tailored exclusively to the scenario of open parking spaces.
@@
In summary, this paper bridges a gap in previous approaches. It empowers drones to make strategic choices in flying directions and recharging locations by predicting traffic flow, while enabling them to autonomously optimize their sensing operations in a decentralized manner.
@@
Notation
Explanation
$u,U,\mathcal{U}$
Index of a drone; total number of drones; set of drones
$m,M,\mathcal{M}$
Index of a charging station; total number of charging stations; set of charging stations
$n,N,\mathcal{N}$
Index of a grid cell; total number of grid cells; set of grid cells
$s,S,\mathcal{S}$
Index of a timeslot; total number of timeslots in a period; set of timeslots
$t,T,\mathcal{T}$
Index of a period; total number of periods; set of periods
$a^{u}$
Period-by-period action taken by $u$
$\mathcal{A}^{u},p^{u}_{n,s}$
Plan of $u$; value of the plan of $u$ at cell $n$ and timeslot $s$
$\mathcal{A},p_{n,s}$
Aggregated plans of all drones; value of the aggregated plan at cell $n$ and timeslot $s$
$\mathcal{R},r_{n,s}$
Target; target value at cell $n$ and timeslot $s$
$V_{n,s}$
Required sensing value at cell $n$ and timeslot $s$
$V^{\prime}_{n,s}$
Overall sensing value collected by all drones at cell $n$ and timeslot $s$
$\hat{V}_{n,s}$
Predicted required sensing value at cell $n$ and timeslot $s$
$G^{a^{u}}$
Group of generated plans when $u$ takes the action $a^{u}$
$l,L$
Index of a plan in $G^{a^{u}}$; total number of plans in $G^{a^{u}}$
$P^{\mathsf{f}}_{u},t^{\mathsf{f}}$
Flying power consumption of $u$; flying time
$P^{\mathsf{h}}_{u},t^{\mathsf{h}}_{u}$
Hovering power consumption of $u$; hovering time
$J^{u},K^{a^{u}}$
Number of visited cells (mobility range) of $u$; range for $u$ selecting cells to visit
$\beta$
Behavior of an agent in planning optimization
$c^{u},e^{u}$
Current location of $u$; current energy consumption of $u$
$o^{u}_{t},r^{u}_{t}$
Observation of $u$ at period $t$; reward of $u$ at the period $t$
$k,H$
Index of a sampled transition; number of sampled transitions
$\gamma,\epsilon$
Discount factor; clip interval hyperparameter
$Q(\cdot),\theta^{\pi}$
Critic network; parameter of critic network
$\pi(\cdot),\theta^{\pi}$
Actor network; parameter of actor network
$\omega_{n,s},\overline{v}$
Prediction coefficient to calculate $\hat{V}_{n,s}$; threshold to remove the regions with low sensing requirements
$I,|S|,|A|$
Total number of iterations in EPOS; number of states in DRL; number of actions in DRL
III Framework Overview
@@ III-A Scenario and Modeling
Matrix of plans. To explain the short-term navigation and sensing of drones over the cells and timeslots in a period $t$, $t\in\mathcal{T}$, the plans of drones $\{\mathcal{A}^{1}(t),\mathcal{A}^{2}(t),...,\mathcal{A}^{U}(t)\}$ are introduced. Each plan is encoded by a matrix of size $N\times S$, with each element represented as $p^{u}_{n,s}(t)\in\{0,1\}$. Here, $p^{u}_{n,s}(t)=1$ denotes that the drone $u$ hovers and collects all required data at cell $n$ at timeslot $s$, whereas for $p^{u}_{n,s}(t)=0$ the drone does not hover at that cell at that time.
@@
Action. To explain the long-term navigation and sensing over all periods, we introduce the period-by-period actions of each drone $a^{u}=\{0,1,2,...,8\}$, meaning to control $u$ to move horizontally along eight directions, which are 1 = north (N), 2 = east (E), 3 = south (S), 4 = west (W), 5 = northeast (NE), 6 = southeast (SE), 7 = southwest (SW), and 8 = northwest (NW), or return to the origin ($a^{u}=0$). Under each action, $u$ executes a plan $\mathcal{A}^{u}(t)$. After completing its sensing tasks, $u$ flies back to one of the charging stations to recharge fully and resume work in the next period $t+1$.
Matrix of aggregated plan. The matrix of the aggregated plan is defined as $\mathcal{A}(t)=\{p_{n,s}(t)\}$, where the element $p_{n,s}(t)$ denotes whether the cell $n$ in the map is covered by drones at timeslot $s$ at period $t$. Note that when multiple drones cover the same cell simultaneously, they collect duplicate sensor data, causing sensing inefficiency. Thus, we treat them as a single entity during over-sensing situations, i.e., $p_{n,s}(t)=min(\sum^{U}_{u=1}p^{u}_{n,s}(t),1)$.
@@
Matrices of required and collected data. In the context of a sensing task, each cell at a timeslot has specific sensing requirements that determine the data acquisition goal of drones. Such sensing requirements can be determined by city authorities as a continuous kernel density estimation, for example, monitoring cycling risk based on requirements calculated by past bike accident data and other information [16, 32]. The high risk level of a cell at a timeslot represents the high importance of sensing (e.g., the crucial intersection of traffic flow), and thus a high number of required sensing values is set. The matrix of required sensing values is denoted as $V(t)=\{V_{n,s}(t)\}$, and the matrix of overall actual sensor data collected by all drones as $V^{\prime}(t)=\{V^{\prime}_{n,s}(t)\}$, where $V_{n,s}(t)$ and $V^{\prime}_{n,s}(t)$ denote the required and collected value at cell $n$ and timeslot $s$ respectively. Note that drones collect all required data of a cell within a timeslot, and thus $V^{\prime}(t)=\mathcal{A}(t)\circ V(t)$, where $\circ$ represents the element-wise product of the two matrices.
@@
Matrix of target. In a real-world scenario, drones lack prior knowledge of the amount of required data values before they begin their sensing operations. Therefore, it becomes essential to build a target on the fly that instructs the drones regarding when and where they should or should not fly to given information from the environment. The matrix of the target is defined as $\mathcal{R}(t)$, which denotes the sensing requirements for all drones at period $t$. The element of the target is denoted as $r_{n,s}(t)=\{0,1\}$, $n\in\mathcal{N}$, $s\in\mathcal{S}$, where $r_{n,s}(t)=1$ requires only a drone to visit the cell $n$ at timeslot $s$, and $r_{n,s}(t)=0$ does not require sensor data collection by a drone.
@@
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a unique altitude while moving between cells to prevent collisions [33]. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.

### v3

@@ Abstract
Swarms of autonomous interactive drones, with the support of recharging technology, can provide compelling sensing capabilities in Smart Cities, such as traffic monitoring and disaster response. This paper aims to deliver a novel coordination solution for the cost-effective navigation, sensing, and recharging of drones. Existing approaches, such as deep reinforcement learning (DRL), offer long-term adaptability, but lack energy efficiency, resilience, and flexibility in dynamic environments. Therefore, this paper proposes a novel approach where each drone independently determines its flying direction and recharging place using DRL, while adapting navigation and sensing through distributed optimization, which improves energy-efficiency during sensing tasks. Furthermore, drones efficiently exchange information while retaining decision-making autonomy via a structured tree communication model. Extensive experimentation with datasets generated from realistic urban mobility underscores an outstanding performance of the proposed solution compared to state-of-the-art methods. Significant new insights show that long-term methods optimize scarce drone resource for traffic management, while the integration of short-term methods is crucial for advising on charging policies and maintaining battery safety.
@@ I Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities [1]. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as reporting traffic congestion at early stage [2], and mapping natural disasters [3]. However, the inherent limitations in battery capacity of drones influence their spatio-temporal coverage. On the other hand, it requires a high cost to build and maintain a large-scale drone swarm together. Therefore, efficient resource management is essential for long operations of drones by ensuring optimal drone recycling and preventing battery depletion. Recently, scholars have introduced the recharging technologies and commercially wireless charging stations for long-term sensing [4].
@@
Inspired by these studies, this work tackles the problem of coordinating swarms of drones for improved distributed navigation, sensing and recharging. The goal is to enhance sensing quality in a large-scale area using both drones and charging stations while minimizing the overall involved costs. This endeavor requires highly adaptive and optimized drone operations. Previous work [5] employs distributed optimization methods for efficient drones path planning and task assignment to enhance sensing quality. These methods enable drones to autonomously self-organize and self-assign their tasks collectively, while respecting privacy and autonomy of drones. However, these methods are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically. The navigation and sensing of drones are significantly influenced by two factors: their flying directions and future requirements for sensor data on the map. For example, drones equipped with insights about the expected increase in traffic flow in certain areas can proactively fly to those locations, even if vehicle density is currently low. This strategic decision can be more effective to allocate limited drone resources for tasks like vehicle detection, although in short-term the benefits are not evident. This underscores the "slower is faster" effect, highlighting the need for long-term planning to optimize overall effectiveness and efficiency in drone sensing operations.
@@
Deep reinforcement learning (DRL) tackles the challenge of long-term optimization. It accounts for long-term and iterative actions through the Bellman equation, which considers the discounted cumulative reward of agents’ actions [6, 7]. Drones, as heterogeneous agents with distinct behavior patterns and action spaces, can benefit from the powerful deep neural networks employed in DRL, enabling an effective handling of sophisticated state space and time-varying environments. Nevertheless, as the number of drones and time periods increase, the curse of dimensionality constrains efficient sensing. For instance, considering a scenario where drones are tasked with learning optimal path planning over a broad range of time scales using DRL. In this setting, the high number of epochs may inadvertently compel the drones to take unnecessary actions. As a result, drones could choose longer paths, leading to higher energy consumption. Furthermore, current methods overlook the need for drones to become more flexible and adaptive to real-time environmental changes while maintaining system resilience, against single points of failure. Addressing these limitations requires reducing actions in a temporal scale as well as effective autonomous mission planning distributively, but no existing approaches incorporate these novelties.
@@
Therefore, in this paper, we propose a new approach for the coordination of swarms of drones, Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a novel multi-agent DRL algorithm that builds upon on a collective learning approach  [8, 9]. DO-RL relies on a learning model to determine the drones’ strategic navigation and flying directions for sensing, which also determines the recharging stations, while leaving the self-planning and selection of specific sensing operations to distributed optimization, i.e., collective learning. This approach significantly reduces the number of decisions in a temporal scale, avoiding unnecessary energy consumption. Then, we design a structured tree communication model within which each agent interacts with its children and parent in a bottom-up and top-down fashion, obtains the aggregated observation to calculate immediate reward in DRL, and independently decides its navigation and sensing. This ensures scalability (support of a large number of agents), efficiency (low communication and computational cost), decentralization and resilience [10, 11, 12]. Furthermore, a set of auxiliary approaches is introduced, including a plan generation strategy that provides high-quality and diverse options for drones navigation and sensing, a plan selection, a periodic state update and a centralized training and decentralized execution framework. To evaluate the effectiveness of DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
@@
The contributions of this paper are outlined as follows: (i) The first attempt to coordinate navigation, sensing, and recharging by a swarm of autonomous and self-planning drones in a distributed and cost-effective way. (ii) A novel approach, DO-RL, to address this problem using both multi-agent DRL and distributed optimization. (iii) The applicability of a designed structured tree communication model for efficient planning and information sharing among agents to ensure decentralization, resilience, and flexibility. (iv) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (v) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (vi) An open-source implementation of the proposed approach and an open dataset [13] containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
@@ II Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as a combinatorial optimization problem, resembling the Traveling Salesmen Problem [5]. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The digraph-based methods, such as greedy approach, [21], ant colony optimization [22], and genetic algorithm [23], are introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, inspection delay, events detection rate and the cost of flying trajectories. However, these works lack comprehensive evaluation in complex sensing scenarios, such as varying time periods, charging stations, and different sensing distributions. Moreover, these centralized methods suffer from a risk of single points of failure [24], wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms.
@@
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective when addressing the distributed sensing problem with drones using complex optimization objectives, limited energy consumption, and a large number of heterogeneous agents. Afrin et al. [26] designed a DRL-based model for edge servers to dynamically allocate tasks to robots (e.g., drones), aimed at enhancing system resilience in smart farming environments. Ding et al. [17] integrated the mobile crowdsensing (i.e., human equipped with mobile sensing devices such as smartphones) into UAV sensing, which overcomes the limitations of both human mobility and drone battery capacity. Omoniwa et al. [18] presented a decentralized approach that leverages DRL and shared information from neighbors to improve the energy efficiency of drones in providing wireless connectivity to ground users in dynamic environments. Nevertheless, these approaches lack privacy and autonomy for drones in self-assignment tasks, reducing flexibility, adaptability to the real-time environmental changes and increasing vulnerability to single points of failure. Moreover, they are confined to a small number of drones and time periods due to the huge action space. In contrast, our approach empowers agents to determine overall flying directions every time period, leaving specific navigation and sensing tasks in each timeslot to planning and selection. This strategic allocation significantly filters out unnecessary actions and reduces the corresponding energy consumption.
Recharging assignment. Ealier work [27, 28] has studied the recharging scheduling problem: drones efficiently collect sensor data while charging from stations. However, scalability becomes a challenge since the algorithms used in these works only deal with a small number of drones. Zhao et al. [19] proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces, specifically for detecting available parking spots. Their solution incorporates the task selection and scheduling using DRL, and adaptive charging scheduling. However, this approach is centralized and tailored exclusively to the scenario of open parking spaces.
@@
In summary, this paper bridges a gap in previous approaches. It empowers drones to make strategic choices in flying directions and recharging locations by predicting traffic flow, while enabling them to autonomously optimize their sensing operations, ensuring decentralization, resilience, and flexibility.
@@
Notation Explanation
$u,U,\mathcal{U}$ Index of a drone; total number of drones; set of drones
$m,M,\mathcal{M}$ Index of a charging station; total number of charging stations; set of charging stations
$n,N,\mathcal{N}$ Index of a grid cell; total number of grid cells; set of grid cells
$s,S,\mathcal{S}$ Index of a timeslot; total number of timeslots in a period; set of timeslots
$t,T,\mathcal{T}$ Index of a period; total number of periods; set of periods
$a^{u}$ The action or flying direction taken by $u$
$P^{u},p^{u}_{n,s}$ The plan of $u$; value of $P^{u}$ at cell $n$ and timeslot $s$
$P^{-u},p^{-u}_{n,s}$ The observed plans of drones by $u$ excluding $P^{u}$; value of $P^{-u}$ at cell $n$ and timeslot $s$
$p_{n,s}$ The aggregated plans of all drones at cell $n$ and timeslot $s$
$V_{n,s}$ Required sensing value at cell $n$ and timeslot $s$
$v_{n,s}$ Collected sensing value at cell $n$ and timeslot $s$
$\mathcal{R},r_{n,s}$ Target; target value at cell $n$ and timeslot $s$
$l,L$ Index of a plan in $G^{a^{u}}$; total number of plans in $G^{a^{u}}$
$C^{\mathsf{f}}(u),t^{\mathsf{f}}$ Flying power consumption of $u$; flying time
$C^{\mathsf{h}}(u),t^{\mathsf{h}}_{u}$ Hovering power consumption of $u$; hovering time
$c^{u},e^{u}$ Current location of $u$; current energy consumption of $u$
$K(a^{u})$ Indexes of cells within the searching range along flying direction $a^{u}$
$J(a^{u})$ Indexes of visited cells within $K(a^{u})$
$\beta$ Behavior of an agent in planning optimization
$o^{u}_{t},r^{u}_{t}$ Observation of $u$ at period $t$; reward of $u$ at the period $t$
$k,H$ Index of a sampled transition; number of sampled transitions
$\gamma,\epsilon$ Discount factor; clip interval hyperparameter
$Q(\cdot),\theta^{Q}$ Critic network; parameter of critic network
$\pi(\cdot),\theta^{\pi}$ Actor network; parameter of actor network
$\omega_{n,s},\overline{v}$ Prediction coefficient to calculate $\hat{V}_{n,s}$; threshold to remove the regions with low sensing requirements
$I,\mathcal{X},A$ Total number of iterations in EPOS; size of state space; number of actions in DRL
$\mathcal{E},W$ Number of episodes; number of neutrals per layer in DNN
III System Model
@@ III-A Scenario and Assumptions
Action. To explain the long-term navigation and sensing over all periods, we introduce the period-by-period actions of each drone $a^{u}=\{0,1,2,...,8\}$, $u\in\mathcal{U}$, meaning to control $u$ to move horizontally along eight directions, which are 1 = north (N), 2 = east (E), 3 = south (S), 4 = west (W), 5 = northeast (NE), 6 = southeast (SE), 7 = southwest (SW), and 8 = northwest (NW), or return to the origin ($a^{u}=0$). Under each action, $u$ executes a short-term navigation and sensing. After completing its sensing tasks, $u$ flies back to one of the charging stations to recharge fully and resume work in the next period $t+1$.
Matrix of plans. To explain the short-term navigation and sensing of drones over the cells and timeslots in a period $t$, $t\in\mathcal{T}$, the plan of a drone $u$ that travels through the direction of $a^{u}$ is introduced, denoted by $P^{u}(a^{u},t)$. It represents the specific navigation and sensing details, including the visited cells and corresponding energy consumption. The plan $P^{u}(a^{u},t)$ is encoded by a matrix of size $N\times S$, with each element represented as $p^{u}_{n,s}(a^{u},t)\in\{0,1\}$. Here, $p^{u}_{n,s}(a^{u},t)=1$ denotes that the drone $u$ hovers and collects all required data at cell $n$ at timeslot $s$, whereas for $p^{u}_{n,s}(a^{u},t)=0$ the drone does not hover at that cell at that time. Moreover, $P^{-u}(a^{u},t)=\{p^{-u}_{n,s}(a^{u},t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$ denotes the observed plans by drone $u$ in time period $t$, indicating that it incorporates and sums the plans of all other drones excluding its own. We formulate the aggregated plans of all drones observed by $u$ at cell $n$ and timeslot $s$ as follows:
$p_{n,s}(a^{u},t)=p^{u}_{n,s}(a^{u},t)+p^{-u}_{n,s}(a^{u},t).$ (1)
@@
Matrix of required sensing values. In the context of a sensing task, each cell at a timeslot has specific sensing requirements that determine the data acquisition goal of drones. Such sensing requirements can be determined by city authorities as a continuous kernel density estimation, for example, monitoring cycling risk based on requirements calculated by past bike accident data and other information [16, 31]. The high risk level of a cell at a timeslot represents the high importance of sensing (e.g., the crucial intersection of traffic flow), and thus a high number of required sensing values is set. The matrix of required sensing values is denoted as $V(t)=\{V_{n,s}(t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$, where $V_{n,s}(t)$ denotes the required sensing value at cell $n$ and timeslot $s$ in the time period $t$. Based on actual sensing performance of drones according to Eq.(1), the sensing value collected by all drones observed by $u$ at cell $n$ and timeslot $s$ is formulated as:
$v_{n,s}(a^{u},t)=p_{n,s}(a^{u},t)\cdot V_{n,s}(t).$ (2)
@@
Matrix of target. In a real-world scenario, drones lack prior knowledge of the amount of required sensing values before they begin their sensing operations. Therefore, it becomes essential to build a target on the fly that instructs the drones regarding when and where they should or should not fly to given information from the environment. The matrix of the target is defined as $\mathcal{R}(t)$, which denotes the sensing requirements for all drones at period $t$. The element of the target is denoted as $r_{n,s}(t)=\{0,1\}$, $n\in\mathcal{N}$, $s\in\mathcal{S}$, where $r_{n,s}(t)=1$ requires only a drone to visit the cell $n$ at timeslot $s$, and $r_{n,s}(t)=0$ does not require sensor data collection by a drone.
@@
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a unique altitude while moving between cells to prevent collisions [32]. We also fix the ground speed of drones and the wind speed for simply calculating power consumption. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. We assume that each charging period is of equal duration and provides sufficient time for the drones to be fully charged. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.